x-4y= 28
-2x-5y=-48

Answer by Alan3354(69443)

x-4y= 28
-2x-5y=-48
=============
2x- 8y = 56 ... 1st eqn times 2
-2x-5y =-48
------------ Add
-13y = 8
y = -8/13
==================
x-4y= 28
x + 32/13 = 28
x = 332/13
==================